OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of 1/2? 2x + y = 4 2x – y = 4 x + 2y = 4 x – 2y = 4

OpenStudy (anonymous):

are these the multiple choice answers given?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

hmm, if you put this into the equation y-y1 = m(x-x1)... the slope is never 1 or 2... but 1/2 as the problem says. thus y - (-3) = 1/2(x-(-2)) .. perhaps im wrong..

OpenStudy (anonymous):

By 1/2 I meant one half if that's what the problem is...

OpenStudy (anonymous):

well if u plug in 1/2 as m (slope) and -2 as x1 and -3 as x2.. u get y = 1/2x - 2 for the equation of a line passing through (-2,-3) with a slope of 1/2

OpenStudy (anonymous):

-3 as y1 i mean

OpenStudy (anonymous):

x-2y=4 =-2y = -x +4 / -2 = y = 1/2x - 2 so x-2y = 4 must be it.